General Announcement: - Tobias Clever informs on the gas maintenance system: the central gas supply is currently disconnected; One shall not bypass it by installing other gas bottles as the gas monitoring alarm is not working and thus pose a safety threat if leakage happens. By the end of this week the activity is completed and possibly by mid of the week gas may be available again. CTRL: - Proceeding with tests of Karabo3; some instruments will be contacted this week for tests. DA: - All fine. EEE: - Nothing to report. VAC: - Nothing to report. XPD: - Nothing to report. DET: - Question on gases activity: XHQ lab areas are also affected, including clean rooms. ITDM: - E-mail will be sent later informing on next week's (Tuesday) interruption of DAQ storage due to a maintenance. XRO: - Not present. LAS: - Timing issues were happening but solved with help from EEE. UNSYS: - Nothing to report. SPB: - Nothing to report. FXE: - Nothing to report. MID: - All fine; no issues. HED: - Nothing to report. SCS: - Nothing to report. SQS: - Nothing to report. SXP: - Nothing to report.